What is: 5/6(4 1/5 * 3/7 - 1/2) + 1/2?
Answer by rfer(16322)

5/6(21/5*3/7-1/2)+1/2
5/6(63/35-1/2)+1/2
5/6(126/70-35/70+35/70)+1/2
5/6*63/35+1/2
315/210+105/210=410/210=1 20/21